You have a string with a mass of 0.0139 kg. You stretch the string with a force of 9.59 N, giving it a length of 1.85 m. Then you vibrate the string transversely at precisely the frequency that corresponds to its fourth normal mode, that is, at its fourth harmonic. What is the wavelength of the standing wave you create in the string? What is the frequency?
Answer #1

a) lambda = 2 L / 4 = 2 * 1.85 / 4

wavelength = 0.925 m

b) v = sqrt [F / (m/l)] = sqrt [9.59 / (0.0139/1.85)] = 35.73 m/s

frequency = v / lambda = 35.73 / 0.925

frequency = 38.62 Hz
Answer #2

Wavelength

For a standing transverse wave in a stretched string that is attached at both ends, the number of the normal mode equals the number of half wavelengths that are contained in its length. The wavelength λn of the nth normal mode is therefore

λn=2Lnn=1,2,3,...

where L denotes the string's length.

In the question, you are given L and n. Substitute to obtain the numerical answer for the wavelength.

λ4=2(1.91 m)4=0.955 m

Frequency

The frequency of the standing wave in the nth normal mode, which is the nth harmonic fn, can be found using the relation that is valid for all periodic waves

fn=vλn

where λn is the wavelength of the nth normal mode (see the formula for wavelength) and v is the speed of transverse waves in the string. This speed is not given directly, but can be obtained from the formula

v=Fm/L

where F is the given tension in the string and m and L are its given mass and length, respectively. Perform all necessary substitutions to obtain the frequency f4 in terms of given quantities.

fn=n2LFm/Ln=1,2,3,...

For the value of the fourth harmonic,




answered by: Muhammad Aslam